Derive integrated law for first order reactions
1
Expert's answer
2020-02-21T07:50:30-0500

Solution.

"A\to B"

"\frac{d[B]}{dt} = -\frac{d[A]}{dt} = k \times [A]"

"-\frac{d[A]}{[A]} = kdt"

Integrate:

"-ln(\frac{[A]}{[A0]}) = k \times t"

"ln(\frac{[A0]}{[A]}) = k \times t"

Answer:

"ln(\frac{[A0]}{[A]}) = k \times t"